A sample of nitrogen at 20 o C was compressed from 300 mL to 0.360 mL and its
new pressure was found to be 400.0 kPa. What was the original pressure in
kPa?
1
Expert's answer
2020-05-19T08:44:59-0400

pV=nRT; n=pV/RT = 400 x 0.36 / 8.314 x 293 = 0.06.

p= 0.06 x 8.314 x 293 / 300 = 0.48 kpa.